OpenStudy (anonymous):

Need help finding answer to sin(2sin^-1(x)) in terms of x.

OpenStudy (anonymous):

\[\sin(2\theta)=2\sin(\theta)\cos(\theta)\]

OpenStudy (anonymous):

here \(\theta=\sin^{-1}(x)\) making \[\sin(\theta)=\sin(\sin^{-1}(x))=x\]

OpenStudy (anonymous):

and \[\cos(\theta)=\cos(\sin^{-1}(x))=\sqrt{1-x^2}\]

OpenStudy (anonymous):

multiply them together, stick a 2 in front of them and you are done

OpenStudy (anonymous):

So that would be 2xsqrt(1-x^2)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you so much :)